A gas in a cylinder fitted with a piston undergoes a cycle composed of three processes. First the gas expands at constant pressure with a heat addition of 404 J and a work output of 10.0 J. Then it is cooled at constant volume by a removal of 50.0 J of heat. Finally, an adiabatic process restores the gas to its initial state. Determine(i) the work of the adiabatic process and (ii) the stored energy of the gas at each of the other two states if its stored energy in the initial state is assigned the value of zero. Sketch the cycle on the pressure – volume and temperature – entropy diagrams. 


1
Expert's answer
2021-04-19T17:09:58-0400

1)

"A=Q_1-Q_2=404-50=354~J,"

2)

"\Delta U=Q_3=10~J."